A Carnot heat engine receives heat from a reservoir at 900°C at a rate of 800 kJ/min and rejects the waste heat to the ambient air at 27°C. The entire work output of the heat engine is used to drive a refrigerator that removes heat from the refrigerated space at 25°C and transfers it to the same ambient air at 27°C. Determine (a) the maximum rate of heat removal from the refrigerated space and (b) the total rate of heat rejection to the ambient air.

1
Expert's answer
2021-05-24T15:56:03-0400

Heat engine

"T_L=27\u00b0c=27+273=300K"

"T_H=900\u00b0c=900+273=1173K"

"Q_H=800KJ\/min"


The power output is obtained from the different efficiency relation

"W=Q_H(1-\frac{T_L}{TH})"

Put value


"W=800(1-\frac{300}{1173})KJ/min"

W=595kJ/min 

The rate of heat rejected to the ambient air "Q_L=Q_H-W"

Put value

"Q_L=800-595=205 KJ\/min"

Refrigerator 

"W=595kJ\/min"

"T_H=27\u00b0c=300K"

"T_L=25\u00b0c=273+25=298k"

The rate of heat removed from the refrigerator is obtained by manipulating the different efficiency relation

"Q_L=\frac{W}{\frac{T_H}{T_L}-1,}"

Put values

"Q_L=\frac{595}{\frac{300}{298}-1}" =88655KJ/min

The rate of heat related of the ambient air is

"Q_H=W+Q_L"

"Q_H=(88655 +595)KJ\/min"

"Q_H=89250KJ\/min"

(b)

Total rate of heat rejected to the ambient air is

"Q_{amb}=Q_H+Q_L"

"Q=(89250+205)KJ\/min"

"Q=89455KJ\/min"